\begin{thebibliography}{1}

\bibitem{Chap}
 Chapman, T.A., Topological invariance of whitehead torsion. \textit{Amer. J. Math.}, 96 (3) (1974), 488--497.


\bibitem{EDErdal}
 Dirican~Erdal, E., The adjoint reidemeister torsion for compact 3-manifolds admitting a unique decomposition. \textit{Turk. J. Math.}, 47 (5) (2023), 1469--1480.
 
 \bibitem{Franz}
 Franz, W., {\"U}ber die torsion einer {\"u}berdeckung. \textit{J. Reine. Angew. Math.}, 173 (1935), 245--254.


\bibitem{JLGROSS}
Gross, J., A unique decomposition theorem for 3--manifolds with connected
  boundary. \textit{Trans. Amer. Math. Soc.}, 142 (1969), 191--199.
 
 \bibitem{Porti2}
 Heusener, M., Porti, J., Holomorphic volume forms on representation varieties of
surfaces with boundary. \textit{Ann. Henri Lebesgue.}, 3 (2020), 341--380.

\bibitem{RCLC}
 Kirby, R.C., Siebenmann, L.C., On the triangulation of manifolds and the hauptvermutung. \textit{Bull. Amer. Math. Soc.},
75 (4) (1969), 742--749.

\bibitem{Lescop}
Lescop, C., Global Surgery Formula for the {C}asson-{W}alker Invariant.
 Princeton University Press, Princeton, NJ, 1996.

\bibitem{Milnor2}
 Milnor, J., A duality theorem for reidemeister torsion. \textit{Ann. Math.}, 76 (1) (1962), 137--147.

\bibitem{Milnor}
 Milnor, J., Whitehead torsion. \textit{Bull. Amer. Math. Soc.}, 72 (3) (1966), 358--426.

\bibitem{Porti}
Porti, J., Torsion de reidemeister pour les vari{\'e}t{\'e}s hyperboliques.
\textit{Mem. Amer. Math. Soc.}, 612 (128) (1997).


\bibitem{Porti3}
Porti, J., Yoon, S., The adjoint Reidemeister torsion for the connected sum of knots. \textit{Quantum Topol.}, 14 (3), (2023), 407--428. 
%https://doi.org/10.4171/QT/180

\bibitem{Reidemeister-Franz}
Reidemeister, K., Homotopieringe und linsenraeume. \textit{Abh. Math. Semin. Univ. Hambg.}, 11 (1935), 102--109.

\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{tikz}
\usetikzlibrary{tqft}
\usetikzlibrary{shapes}
\usepackage{cite}
\usepackage{tikz}
\usetikzlibrary{tqft}
\usepackage{tikz-cd} 
\usepackage[margin=1.40in]{geometry}

 \newtheorem{theorem}{Theorem}[subsection]
 
 \newtheorem{lemma}[theorem]{Lemma}

 \theoremstyle{definition}
 \newtheorem{definition}[theorem]{Definition}
 \theoremstyle{remark}
 \numberwithin{equation}{subsection}

\begin{document}

\title{On twisted torsion of compact $3$-manifolds}
\author{ESMA D\.IR\.ICAN ERDAL}

\address[1]{%
Deptartment of Mathematics\\
I\c{s}{\i}k University\\
 \.{I}stanbul\\
  T\"{u}rkiye}
\email{esma.diricanerdal@isikun.edu.tr}

\subjclass[2010]{Primary 55U99; Secondary (optional) 18G99, 57Q10}
\keywords{Reidemeister torsion, connected sum, orientable closed manifold}

\begin{abstract}
	Let $M$ be a $3$-manifold with connected non-vacuos boundary which is not spherical. Assume that $N$ is another $3$-manifold with vacuous boundary and $N^{\ast}$ is the $3$-manifold obtained by removing from $N$ the interior of a $3$-cell. In the present paper, we find a relationship between the multiplicative property of the twisted Reidemeister torsion and the connected sum operation on these manifolds in order to understand their topology and geometry.
\end{abstract}	
\maketitle

\section{Introduction}
\label{Sec:1}
Reidemeister torsion is a piecewise linear topological invariant for $3$-manifolds. It was first defined by Reidemeister and Franz to give a homeomorphism classification of lens spaces \cite{Reidemeister-Franz,Franz}. Kirby and Siebenmann showed that the Reidemeister torsion is a topological invariant for manifolds \cite{RCLC}. The invariance for arbitrary simplicial complexes was proved by Chapman \cite{Chap}, and thus the classification of lens spaces of Reidemeister and Franz was shown to be a topological invariant. Thanks to the Reidemeister torsion, Milnor disproved Hauptvermutung by constructing two homeomorphic but combinatorially distinct finite simplicial complexes. Moreover, he described Reidemeister torsion with the Alexander polynomial which plays an important role in knot theory and links \cite{Milnor2,Milnor}. Reidemeister torsion has proven its utility in a number of topics in three-dimensional topology. For instance, Lescop defined the Casson-Walker-Lescop invariants by using Reidemeister torsion \cite{Lescop}. 

Throughout this paper any $3$-manifold considered here is triangulated oriented connected and compact with connected non-empty boundary and $G$ is considered as a complex reductive algebraic group $SL_n(\mathbb{C})$ or $PSL_n(\mathbb{C})$ with its Lie algebra $\mathfrak{g}$. Let $X$ be a finite CW-complex with the cell-decomposition $K$ and let $\rho: \pi_1(X)\rightarrow G$ be any representation. We denote the twisted chain 
complex by $C_{\ast}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})$ and the twisted $p$-th homology group of $C_{\ast}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})$ by $$H_{p}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})=Z_{p}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})/ B_{p}(K;\mathfrak{g}_{\mathrm{Ad}_\rho}),$$ 
where
$Z_{p}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})=\mathrm{Ker}\{\partial_{p}\otimes \mathrm{id}:C_{p}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})\rightarrow
C_{p-1}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})\}$ is the subspaces of cycles and
 $B_{p}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})=\mathrm{Im}\{\partial_{p+1}\otimes \mathrm{id}:C_{p+1}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})\rightarrow
C_{p}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})\}$ is the subspaces of boundaries. 
For a given homology basis $\mathbf{h}^{M}_p$ of $H_p(M;\mathfrak{g}_{\mathrm{Ad}_{\rho}}),$  $\mathbb{T}_{\rho}(M,\{\mathbf{h}^{M}_p\}_{p=0}^3)$ denotes the twisted Reidemeister torsion of $M$ twisted by the representation $\rho: \pi_1(M)\rightarrow G.$

Finding a relationship between the multiplicative property of the twisted Reidemeister torsion and the connected sum operation on these manifolds is crucial to understand the topology and geometry of manifolds. For $2$-dimensional manifolds, for example the planar surface with four boundary components, Porti gives a formula that computes the twisted Reidemeister torsion of this planar surface by considering this surface as the connected sum of two pairs of pants that intersect in precisely two points \cite{Porti2}. For closed $3$-dimensional manifolds, the twisted Reidemeister torsion for the connected sum of knots is given in \cite{Porti3}. Our aim is to fill this gap in the literature for the twisted Reidemeister torsion of compact $3$-dimensional manifolds. We use the following homeomorphism between the connected sum $M\# N$ and the disk sum $M\bigtriangleup N^{\ast}$ given in \cite{JLGROSS} as follows
 $$M\# N\approx M\bigtriangleup N^{\ast}.$$ By using this homeomorphism together with \cite[Theorem 1.2]{EDErdal}, we obtain the following theorem.

  \begin{theorem}\label{primetheo}
Assume that $M$ is a $3$-manifold with connected non-vacuos boundary which is not spherical and $N$ is another $3$-manifold with vacuos boundary. Assume also that 
$\varrho: \pi_1(M\# N)\rightarrow G$ is a given representation. For a given basis $\mathbf{h}_i^{M\# N}$ of $H_i(M\# N;\mathfrak{g}_{\mathrm{Ad}_{\varrho}})$, there exist homology bases $\mathbf{h}^{M}_i$ and $\mathbf{h}^{N^{\ast}}_i$ of $H_i(M;\mathfrak{g}_{\mathrm{Ad}_{\psi_{_{M}}}})$ and $H_i(N^{\ast};\mathfrak{g}_{\mathrm{Ad}_{\psi_{_{N^{\ast}}}}})$ such that the following formula holds
\begin{eqnarray*}
&\mathbb{T}_\varrho(M\# N,\{\mathbf{h}_i^{M\# N}\}_{p=0}^3) = &\mathbb{T}_{{\psi_{_1}}}(M,\{\mathbf{h}_i^{M}\}_{p=0}^3) \; \mathbb{T}_{{\psi_{_2}}}(N^{\ast},\{\mathbf{h}_i^{N^{\ast}}\}_{p=0}^3).
\end{eqnarray*}
Here,  $\sigma_{|_{\pi_1(M)}}=\psi_{_1},$ $\sigma_{|_{\pi_1(N^{\ast})}}=\psi_{_2}$ are the restrictions of the representation $\sigma=[\varphi^{-1}]\circ\varrho :\pi_1(M\bigtriangleup N^{\ast})\rightarrow G.$
\end{theorem}

The paper is organized as follows. In Section 2, we briefly recall basic definitions on the twisted Reidemeister torsion. We give the details about disk sum and Theorem 1.2 in \cite{EDErdal} in Section 3, and prove Theorem \ref{primetheo} in Section 3.1.


\section{The Twisted Reidemeister Torsion}
 \label{Sec:3}
 Consider an $n$-dimensional CW-complex $X$ with its universal covering 
 $\widetilde{X}.$ Let us denote the non-degenerate Killing form on $\mathfrak{g}$ by $\mathcal{B}$ which is given by $\mathcal{B}(A,B) =  4 \cdot \mathrm{Trace}(AB).$ We consider the action of $\pi_1(X)$ on $\mathfrak{g}$ via the adjoint of $\rho$ for a representation $\rho:\pi_1(X)\rightarrow G.$ 

Let $K$ be a cell-decomposition of $X$ and $\widetilde{K}$ be a lifting of $K.$ 
By using the cellular chain complex $C_{\ast}(\widetilde{K}; \mathbb{Z}),$ the twisted chain complex is defined as follows 
 \begin{equation}\label{defchn1}
C_{\ast}(K;\mathfrak{g}_{\mathrm{Ad}_\rho}):=\displaystyle
C_{\ast}(\widetilde{K};\mathbb{Z})\displaystyle\otimes \mathfrak{g}
/\sim, \end{equation}
where $\sigma\otimes t \sim
\gamma\cdot\sigma\otimes\gamma\cdot t$ for every $\gamma\in \pi_1(X),$
$\pi_1(X)$ acts on $\widetilde{X}$ by deck
transformations, and the action of $\pi_1(X)$ on $\mathfrak{g}$ is
the adjoint action.

Let $\{e_1^p,\ldots, e_{m_p}^p\}$ be the set of $p$-cells of $K$ that provides us a 
$\mathbb{Z}$-basis for $C_p(K;\mathbb{Z}).$ Let us choose a lift $\widetilde{e}^p_j$ of $e^p_j$ for each $j=1,\ldots, m_p.$ Then we obtain a $\mathbb{Z}[\pi_1(X)]$-basis $c_p=\{\widetilde{e^p_j}\}_{j=1}^{m_p}$ of $C_p(\widetilde{K};\mathbb{Z}),$ where $\mathbb{Z}[\pi_1(X)]$ is the integral group ring.
 By using $\mathcal{B}$-orthonormal basis
$\mathcal{A}=\{\mathfrak{a}_k\}_{k=1}^{\dim \mathfrak{g}}$ of $\mathfrak{g}$, one can construct a geometric basis $\mathbf{c}_p=c_p\otimes_{\rho}\mathcal{A}$ for $C_p(K;\mathfrak{g}_{\mathrm{Ad}_
\rho}).$ 

Now let us consider the following chain complex  $C_{\ast}:=C_{\ast}(K;\mathfrak{g}_{\mathrm{Ad}_{\rho}})$
\begin{equation*}\label{chaincomplex}
  \begin{array}{ccc}
(0 \to C_{n}(K;\mathfrak{g}_{\mathrm{Ad}_
\rho}) {\rightarrow}
C_{n-1}(K;\mathfrak{g}_{\mathrm{Ad}_
\rho})\rightarrow \cdots \rightarrow C_{0}(K;\mathfrak{g}_{\mathrm{Ad}_
\rho}){\rightarrow}0).
  \end{array}
 \end{equation*}
 For $p\in \{0,\ldots,n\},$ let $H_p(C_{\ast})=Z_p(C_{\ast})/B_p(C_{\ast})$ be
$p$-th homology group of the twisted chain complex $C_{\ast}(K;\mathfrak{g}_{\mathrm{Ad}_{\rho}})$ with the structure of a $\mathbb{C}$-vector space, where
 $$B_p(C_{\ast})=\mathrm{Im}\{\partial_{p+1}:C_{p+1}\rightarrow C_{p}
\},$$
$$Z_p(C_{\ast})=\mathrm{Ker}\{\partial_{p}:C_{p}\rightarrow
C_{p-1} \}.$$ 
Hence, we get the short exact sequences
\begin{equation}\label{Equation1}
0\longrightarrow Z_p(C_\ast) \stackrel{\imath}{\hookrightarrow} C_p(C_{\ast})
\stackrel{\partial_p}{\longrightarrow} B_{p-1}(C_\ast) \longrightarrow 0,
\end{equation}
\begin{equation}\label{Equation2}
0\longrightarrow B_p(C_\ast) \stackrel{\imath}{\hookrightarrow} Z_p(C_\ast)
\stackrel{\varphi_p}{\longrightarrow} H_p(C_\ast) \longrightarrow 0.
\end{equation}
Here, $\imath$ and $\varphi_p$ denote the inclusion and the natural
projection, respectively.

Let $s_p:B_{p-1}(C_{\ast})\rightarrow C_p(C_{\ast})$ and
$\ell_p:H_p(C_{\ast})\rightarrow Z_p(C_{\ast})$ be respectively sections of
$\partial_p:C_p(C_{\ast}) \rightarrow B_{p-1}(C_{\ast})$ and
$\varphi_p:Z_p(C_{\ast})\rightarrow H_p(C_{\ast}).$ Applying Splitting Lemma for the short exact sequences (\ref{Equation1})
and (\ref{Equation2}), we obtain the equation
\begin{equation}\label{Equation0}
C_p(C_{\ast})=B_{p}(C_{\ast})\oplus \ell_p(H_p(C_{\ast }))\oplus s_p(B_{p-1}(C_{\ast})).
\end{equation}
For the bases $\mathbf{b_p}$ and
$\mathbf{h_p}$ of $B_p(C_\ast)$ and $H_p(C_\ast),$ if we use equation~(\ref{Equation0}) then the disjoint union $\mathbf{b}_p\sqcup \ell_p(\mathbf{h}_p)\sqcup s_p(\mathbf{b}_{p-1})$ becomes a new basis for $C_p(C_{\ast}).$

\begin{definition} Let $\left[\mathbf{b}_p\sqcup \ell_p(\mathbf{h}_p)\sqcup
s_p(\mathbf{b}_{p-1}),\mathbf{c}_p\right]$ be the determinant of
the transition matrix from basis $\mathbf{b}_p\sqcup \ell_p(\mathbf{h}_p)\sqcup
s_p(\mathbf{b}_{p-1})$ to $\mathbf{c}_p$
of $C_p(C_{\ast}).$ Then the twisted Reidemeister torsion of a chain complex $C_{\ast}(K;\mathfrak{g}_{\mathrm{Ad}_
\rho})$ is defined by
 $$\mathbb{T}(C_{\ast}(K;\mathfrak{g}_{\mathrm{Ad}_
\rho}),\{\mathbf{c}_p\}_{p=0}^{n},\{\mathbf{h}_p\}_{p=0}^{n})
 =\prod_{p=0}^n \left[\mathbf{b}_p\sqcup \ell_p(\mathbf{h}_p)\sqcup
s_p(\mathbf{b}_{p-1}), \mathbf{c}_p\right]^{(-1)^{(p+1)}}.$$
Here, the twisted Reidemeister torsion lives in $\mathbb{C}^*/\{\pm 1\}.$
\end{definition}
Note that the twisted Reidemeister torsion of $C_{\ast}(K;\mathfrak{g}_{\mathrm{Ad}_\rho})$ depends on the choice of the homology bases 
$\mathbf{h}_p.$ However, it does not depend on the conjugacy class of $\rho$, the choice of the lifts $\widetilde{e}^p_j$ of the cells $e^p_j,$ and the basis 
$\mathcal{A}$ since $\mathcal{A}$ is orthonormal with respect to $\mathcal{B}$ by \cite[Section 0.2, Remarks a1, a2 in p.10]{Porti}. Moreover, this invariant does not depend on the bases $\mathbf{b}_p,$ the sections $s_p, \ell_p$ due to \cite[Section 3, p.365]{Milnor}.

In \cite{Milnor}, Milnor proved that the twisted Reidemeister torsion of $M$ is invariant under subdivisions. Thus, it defines an invariant of manifolds as follows.
\begin{definition}
Suppose that $M$ is a smooth compact $n$-manifold with a triangulation $K,$ where $n=1,2,3.$ Let $\rho:\pi_1(M)\rightarrow G$ be a given representation and let  
$\{\mathbf{h}_p\}_{p=0}^{n}$ be given homology bases. Then the twisted Reidemeister torsion of $M$ is defined by
$$\mathbb{T}_{\rho}(M,\{\mathbf{h}_p\}_{p=0}^{n})=\mathbb{T}(C_{\ast}(K;\mathfrak{g}_{\mathrm{Ad}_
\rho}),\{\mathbf{c}_p\}_{p=0}^{n},\{\mathbf{h}_p\}_{p=0}^{n}).$$ 
Indeed, the twisted Reidemiester torsion is independent of the triangulation $K$ of $M.$
\end{definition}

The following theorem reviews the multiplicativity property of the twisted Reidemeister torsion and it appears to be a very powerful tool for computing the twisted Reidemeister torsions.
\begin{theorem}[See \cite{Porti}]\label{prt1}
Let $X$ be a compact CW-complex with sub-complexes $X_1,$ $X_2 \subset X$ such that $X=X_1\cup X_2$ and $Y=X_1\cap X_2.$ Assume that $Y_1,\ldots,Y_k$ are the connected components of $Y.$ For the inclusions 
$$Y \overset{i_\nu} {\hookrightarrow} X_\nu \overset{j_\nu} {\hookrightarrow}X,$$ let $\rho:\pi_1(X)\rightarrow G$ be a representation with the restrictions 
$\rho_{_{|_{X_{\nu}}}}:\pi_1(X_\nu)\rightarrow G$ and
$\rho_{_{|_{Y_{\mu}}}}:\pi_1(Y_\mu)\rightarrow G,$ $\nu= 1,2,$ $\mu=1,\ldots,k.$ Then the followings hold.
\begin{itemize}
\item[(i)]{The following sequence is short-exact 
\begin{eqnarray*}
&& 0 \rightarrow \underset {{\mu}}{\oplus} 
C_{\ast}(Y_{\mu};\mathfrak{g}_{\mathrm{Ad}_{\rho_{_{|_{Y_{\mu}}}}}})
\overset{(i_1)_{\#} \oplus (i_2)_{\#} }{\longrightarrow }C_{\ast}(X_1;\mathfrak{g}_{\mathrm{Ad}_{\rho_{_{|_{X_{1}}}}}})\oplus C_{\ast}(X_2;\mathfrak{g}_{\mathrm{Ad}_{\rho_{_{|_{X_{2}}}}}})\\
&&\quad \quad\quad \quad\quad\quad\quad \;\; \tikz\draw[->,rounded corners,nodes={asymmetrical rectangle}](6.30,0.4)--(6.30,0)--(1,0)--(1,-0.4)node[yshift=4.0ex,xshift=12.0ex] {${(j_1)_{\#} -(j_2)_{\#}}$};
 \nonumber\\ 
&&\quad\quad\quad\quad\quad C_{\ast}(X;\mathfrak{g}_{\mathrm{Ad}_\rho})\rightarrow 0.
  \end{eqnarray*}}
\item[(ii)]{Corresponding to the sequence in (i), there is a Mayer-Vietoris long exact sequence in homology with twisted coefficients 
\begin{eqnarray}\label{es12es}
& &\mathcal{H}_{\ast}:\cdots  \longrightarrow \underset {{\mu}}{\oplus} 
H_i(Y_{\mu};\mathfrak{g}_{\mathrm{Ad}_{\rho_{_{|_{Y_{\mu}}}}}})\overset{{i}^\ast_1 \oplus {i}^\ast_2 }{\longrightarrow }H_i(X_1;\mathfrak{g}_{\mathrm{Ad}_{\rho_{_{|_{X_{1}}}}}})\oplus H_i(X_2;\mathfrak{g}_{\mathrm{Ad}_{\rho_{_{|_{X_{2}}}}}}) \nonumber \\ 
&&\quad \quad\quad\quad\quad\quad\quad \quad\; \tikz\draw[->,rounded corners,nodes={asymmetrical rectangle}](6.40,0.4)--(6.40,0)--(1,0)--(1,-0.4)node[yshift=4.0ex,xshift=12.0ex] {${j}^\ast_1 - {j}^\ast_2$ };
 \nonumber\\ 
&&\quad\quad\quad\quad\quad  H_i(X;\mathfrak{g}_{\mathrm{Ad}_\rho})\longrightarrow  \underset {\mu}{\oplus} H_{i-1}(Y_{\mu};\mathfrak{g}_{\mathrm{Ad}_{\rho_{_{|_{Y_{\mu}}}}}})\longrightarrow  \cdots
 \end{eqnarray}}
\item[(iii)]{Assume that $\mathbf{h}^X_i$, $\mathbf{h}_i^{X_1}$,$\mathbf{h}^{X_2}_i,$ $\mathbf{h}^{Y_{\mu}}_i$ are respectively basis for $H_i(X;\mathfrak{g}_{\mathrm{Ad}_\rho}),$ $H_i(X_1;\mathfrak{g}_{\mathrm{Ad}_{\rho_{_{|_{X_{1}}}}}}),$ $H_i(X_2;\mathfrak{g}_{\mathrm{Ad}_{\rho_{_{|_{X_{2}}}}}}),$ and $H_i(Y_{\mu};\mathfrak{g}_{\mathrm{Ad}_{\rho_{_{|_{Y_{\mu}}}}}}).$ Then the long exact sequence (\ref{es12es}) can be considered as a chain complex and the following formula is valid
\begin{eqnarray*}
&& \mathbb{T}_{\rho_{_{|_{X_{1}}}}}(X_1,\{\mathbf{h}^{X_1}_i\})\; 
 \mathbb{T}_{\rho_{_{|_{X_{2}}}}}(X_2,\{\mathbf{h}^{X_2}_i\})=\mathbb{T}_{\rho}(X,\{\mathbf{h}^X_i\})
\prod_{\mu=1}^k\mathbb{T}_{\rho_{_{|_{Y_{\mu}}}}}(Y_{\mu},\{\mathbf{h}^{Y_{\mu}}_i\})\\
&& \quad \quad \quad \quad \quad \quad  \quad\quad \quad \quad \quad  \quad   \quad \quad \quad \quad  \; \;  \times \; \mathbb{T}(\mathcal{H}_{\ast}, \{\mathbf{h}_{\ast \ast}\}).
\end{eqnarray*}}
\end{itemize}
\end{theorem} 
 We refer to \cite[Theorem 3.2]{Milnor} and
\cite[Section 0.4, Proposition 0.11]{Porti} for the proof of Theorem \ref{prt1}. 

Note that the torsion $\mathbb{T}(\mathcal{H}_{\ast},\{\mathbf{h}_{\ast \ast}\})$ is called the corrective term and this term can be computed by using the following useful lemma.
\begin{lemma}[See \cite{EDErdal}]\label{crorectivetermlem}
Let $X$ be an $n$-dimensional CW-complex. Then the corrective term satisfies the following formula
\begin{equation*}
\mathbb{T}\left( \mathcal{H}_{\ast},\{\mathbf{h}_p\}_{p=0}^{3n+2},\{0\}_{p=0}^{3n+2}\right)
 =\prod_{p=0}^{3n+2} \left[\mathbf{h}'_p, \mathbf{h}_p\right]^{(-1)^{(p+1)}}.
\end{equation*}
Here, $\mathbf{h}'_p=\mathbf{b}_p\sqcup s_p(\mathbf{b}_{p-1}).$ 
\end{lemma}
\section{Main Result}
 Let $M$ be a $3$-manifold with connected non-vacuos boundary which is not spherical and let $N$ be another $3$-manifold with vacuous boundary. Consider the disk sum $M\bigtriangleup N^{\ast}$ of these manifolds, where $N^{\ast}$ is the $3$-manifold obtained by removing from $N$ the interior of a $3$-cell. The disk sum $M\bigtriangleup N^{\ast}$ can be formed by pasting a $2$-cell on the boundary of $M$ to a $2$-cell on the boundary of $N^{\ast}.$ The operation of disk sum $\bigtriangleup$ is well-defined, associative, and commutative up to homeomorphism. By \cite{JLGROSS}, the manifold $M$ admits a unique decomposition into $\triangle$-prime $3$-manifolds. Using this decomposition, the following theorem shows that the twisted Reidemeister torsion has a multiplicative property on the disk sum decomposition of compact $3$-manifolds.
\begin{theorem}[See Theorem 1.2 in \cite{EDErdal}]\label{theo1}
Consider the disk sum $M=\overset{n}{\underset {i=1}{\bigtriangleup}}(M_i)$ and denote by the disks on the boundaries of $M_j$ and $M_{j+1}$ that are identified in the construction of $M$ by $\mathbb{D}_j^2$ for each $j \in \{1,\ldots,n-1\}.$ Suppose that interiors of $\mathbb{D}_j^2$ 's are pairwise disjoint in $M.$ 
Suppose also that $\varrho: \pi_1(M)\rightarrow G$ is a given representation with the restrictions $\varrho_{|_{\pi_1(M_i)}}=\psi_{_i}.$ For a given basis 
$\mathbf{h}^{M}_p$ of $H_p(M;\mathfrak{g}_{\mathrm{Ad}_{\varrho}})$ and the basis 
$\mathbf{h}_{0}^{\mathbb{D}_j^2}=f^j_{\ast}(\varphi_0(\mathbf{c}_0))$  of $H_0(\mathbb{D}_j^2;\mathfrak{g}_{\mathrm{Ad}_{\varrho_{|_{\mathbb{D}_j^2}}}}),$ there exists a basis 
$\mathbf{h}^{M_i}_p$ of $H_p(M_i;\mathfrak{g}_{\mathrm{Ad}_{\psi_{_{M_i}}}})$  for each $i\in\{1,\ldots,n\}$ such that the following formula is valid
\begin{eqnarray*}
 \mathbb{T}_{\varrho}(M,\{\mathbf{h}^{M}_p\}_{p=0}^3)= \prod_{i=1}^{n}
 \mathbb{T}_{{\psi_{_i}}}(M_i,\{\mathbf{h}_p^{M_i}\}_{p=0}^{3}),
 \end{eqnarray*}
where $f^j_{\ast}$ denotes the map induced by the simple homotopy equivalence $f^j:\{pt.\}\rightarrow \mathbb{D}_j^2$ between a point $pt.$ and disk 
$\mathbb{D}_j^2,$ the map $\varphi_0:Z_0(\{pt.\};\mathrm{Ad}_{\varphi})\rightarrow H_0(\{pt.\};\mathrm{Ad}_{\varphi})$ is the natural projection, and $\mathbf{c}^j_0$ is the geometric basis of $C_{0}(\{pt.\};{\mathrm{Ad}_\varphi}).$ 
\end{theorem} 

By using the homeomorphism $M\# N\approx M\bigtriangleup N^{\ast}$ together with Theorem \ref{theo1}, we show that the twisted Reidemeister torsion has a multiplicative property on the connected sum of compact $3$-manifolds.
\subsection{Proof of Theorem \ref{primetheo}}
 \label{Suec:4}
 
Consider the homeomorphism $\varphi:M\# N \rightarrow M\bigtriangleup N^{\ast}.$ 
At the level of homology there are isomorphisms 
$$[\varphi_i]: H_i(M\# N) \rightarrow H_i(M\bigtriangleup N^{\ast})$$ for each $i=0,1,2,3.$ Using the induced isomorphism $$[\varphi^{-1}]:\pi_1(M\bigtriangleup N^{\ast})\rightarrow \pi_1(M\# N),$$ we get the representation of $\pi_1(M\bigtriangleup N^{\ast})$ as follows
$$\sigma=[\varphi^{-1}]\circ\varrho :\pi_1(M\bigtriangleup N^{\ast})\rightarrow G.$$

 Let $\mathbf{h}_i^{M\# N}$ be a given basis of $H_i(M\# N).$ Since Reidemeister torsion is a topological invariant, we obtain the following formula
\begin{equation}\label{prmequatn4}
\mathbb{T}_\varrho(M\# N,\{(\mathbf{h}_i^{M\# N})\}_{p=0}^3)=
\mathbb{T}_\sigma(M\bigtriangleup N^{\ast},\{[\varphi_i](\mathbf{h}_i^{M\# N})\}_{p=0}^3).
\end{equation}

Recall that the disk sum $M \bigtriangleup N^{\ast}$ can be formed by pasting a topological disk ${\mathbb{D}^2}$ on boundary of $M$ to a topological disk 
${\mathbb{D}^2}$ on boundary of $N^{\ast}.$ Hence, by Theorem~\ref{prt1}, the computation of the twisted Reidemeister torsion of $M \bigtriangleup N^{\ast}$ involves ${\mathbb{T}_{{\sigma_{|_{\mathbb{D}^2}}}}(\mathbb{D}^2,\{\mathbf{h}_0^{\mathbb{D}^2}\})}$. Let us choose the homology basis $\mathbf{h}_{0}^{\mathbb{D}^2}$ of $\mathbb{D}^2$ as $f_{\ast}(\varphi_0(\mathbf{c}_0)),$ where $f_{\ast}$ is the map induced by the simple homotopy equivalence $f:\{pt.\}\rightarrow \mathbb{D}^2,$ $\varphi_0:Z_0(\{pt.\};\mathrm{Ad}_{\varphi})\rightarrow H_0(\{pt.\};\mathrm{Ad}_{\varphi})$ is the natural projection, and $\mathbf{c}_0$ is the geometric basis of $C_{0}(\{pt.\};{\mathrm{Ad}_\varphi}).$ Thus, 
${\mathbb{T}_{{\sigma_{|_{\mathbb{D}^2}}}}(\mathbb{D}^2,\{\mathbf{h}_0^{\mathbb{D}^2}\})}=1$ by 
\cite[Lemma 3.1]{EDErdal}. 

Assume that $\sigma_{|_{\pi_1(M)}}=\psi_{_1}$ and $\sigma_{|_{\pi_1(N^{\ast})}}=\psi_{_2}$ are the restrictions of the representation $\sigma.$ For the basis $[\varphi_i](\mathbf{h}_i^{M\# N})$ of $H_i(M\bigtriangleup N^{\ast})$, by Theorem~\ref{theo1}, there exist homology bases $\mathbf{h}^{M}_i$ and $\mathbf{h}^{N^{\ast}}_i$ of $H_i(M;\mathfrak{g}_{\mathrm{Ad}_{\psi_{_{M}}}})$ and $H_i(N^{\ast};\mathfrak{g}_{\mathrm{Ad}_{\psi_{_{N^{\ast}}}}})$ such that the following formula holds
 \begin{equation}\label{eqnlst}
\mathbb{T}_\sigma(M\bigtriangleup N^{\ast},\{[\varphi_i](\mathbf{h}_i^{M\# N})\}_{p=0}^3) = \mathbb{T}_{{\psi_{_1}}}(M,\{\mathbf{h}_i^{M}\}_{p=0}^3) \; \mathbb{T}_{{\psi_{_2}}}(N^{\ast},\{\mathbf{h}_i^{N^{\ast}}\}_{0}^3).
\end{equation}

Combining equations (\ref{prmequatn4}) and (\ref{eqnlst}) implies that the twisted Reidemeister torsion has a multiplicative property on the connected sum of compact $3$-manifolds without a corrective term.
\begin{eqnarray*}
&\mathbb{T}_\varrho(M\# N,\{\mathbf{h}_i^{M\# N}\}_{0}^3) = &\mathbb{T}_{{\psi_{_1}}}(M,\{\mathbf{h}_i^{M}\}_{p=0}^3) \; \mathbb{T}_{{\psi_{_2}}}(N^{\ast},\{\mathbf{h}_i^{N^{\ast}}\}_{p=0}^3).
\end{eqnarray*}



\end{document}